OpenStudy (anonymous):

solve the following formula for b1 a=(b1+b2/2)h

OpenStudy (anonymous):

ah=(b1+b2/2) ah-b2/2=b1

OpenStudy (tyteen4a03):

This looks like the formula of a trapezoid but you need to remember that it's \(\frac{b1+b2}{2}\), what you typed makes it look like this: \(b1+\frac{b2}{2}\).

OpenStudy (anonymous):

*a/h